# Any function I can actually write down is measurable, right?

### v2

@@ Appendix A An explicit paradoxical decomposition of the sphere
These generate a free group $F\subseteq SO(3)$. This can be seen by analyzing the mod $\mathrm{mod}\ 5$ behavior of the numerators of entries of products of $\rho$ and $\phi$ or by analyzing the dynamics of the matrices in the $5$-adic numbers, as is done in [30]. Let $D$ be the set of points in $S^{2}$ fixed by some element of $F$. Note that since every product of $\rho$ and $\phi$ (and their inverses) is a rational matrix, we have that every element of $D$ is the normalization of a vector with rational coordinates. Therefore for any $\bar{x},\bar{y}\in D$, there is a quartic (i.e., degree $4$) field extension $K$ of $\mathbb{Q}$ such that the coordinates of $\bar{x}$ and $\bar{y}$ are all elements of $K$.
[…]
@@
By A.3 and A.4, we have that the restriction of this equivalence relation to $S^{2}\setminus D$ is $\Delta^{0}_{3}$ (and therefore $\Delta^{1}_{2}$). The result then follows from 3.3. ∎
@@
Let $\tau$ and $\sigma$ be the two generators of $F_{2}$. The four sets $(A_{i})_{i\leq 4}$ given in [31, Thm. 4.2] are defined as follows:
@@
  • $x\in A_{3}$ if and only if the first letter in the reduced word of $x$ is $\sigma^{-1}$ and $x$ is not of the form $\sigma^{-n}$ for some $n$.
@@
This can almost certainly be improved to four pieces, the known optimal size of a paradoxical decomposition of $S^{2}$, by directly adapting the proof of [31, Thm. 4.5]. There is also no fundamental extra subtlety in writing out a $\Pi^{1}_{2}$ decomposition of the full unit ball, rather than just the unit sphere.

### v3

@@ Appendix A An explicit paradoxical decomposition of the sphere
These generate a free group $F\subseteq SO(3)$. This can be seen by analyzing the $\mathrm{mod}\ 5$ behavior of the numerators of entries of products of $\rho$ and $\phi$ or by analyzing the dynamics of the matrices in the $5$-adic numbers, as is done in [34]. Let $D$ be the set of points in $S^{2}$ fixed by some element of $F$. Note that since every product of $\rho$ and $\phi$ (and their inverses) is a rational matrix, we have that every element of $D$ is the normalization of a vector with rational coordinates. Therefore for any $\bar{x},\bar{y}\in D$, there is a quartic (i.e., degree $4$) field extension $K$ of $\mathbb{Q}$ such that the coordinates of $\bar{x}$ and $\bar{y}$ are all elements of $K$.
Now let $\alpha$ be the rotation matrix corresponding to the unit quaternion $\frac{1+i+j\sqrt[3]{2}+k\sqrt[5]{2}}{\sqrt{2+\sqrt[3]{4}+\sqrt[5]{4}}}$. In other words, $\alpha$ represents a rotation of $2\arccos\left(\frac{1}{\sqrt{2+\sqrt[3]{4}+\sqrt[5]{4}}}\right)$ radians about the axis $\langle 1,\sqrt[3]{2},\sqrt[5]{2}\rangle$. By a short field-theoretic argument, [^30] If $K$ is a quartic field extension of $\mathbb{Q}$, then $K(\sqrt[3]{2})$ cannot contain $\sqrt[5]{2}$ by the multiplicativity of degrees of field extensions. the set $\{1,\sqrt[3]{2},\sqrt[5]{2}\}$ is linearly independent over any quartic field extension $K$ of $\mathbb{Q}$. In particular, this implies that for any distinct $\bar{x},\bar{y}\in D$, we have that the inner products of $\bar{x}$ and $\langle 1,\sqrt[3]{2},\sqrt[5]{2}\rangle$ and of $\bar{y}$ and $\langle 1,\sqrt[3]{2},\sqrt[5]{2}\rangle$ are not equal. Since these inner products are preserved by $\alpha$, we have that $\bar{x}\neq\alpha^{n}\cdot\bar{y}$ for any $n$. Moreover, $2\arccos\left(\frac{1}{\sqrt{2+\sqrt[3]{4}+\sqrt[5]{4}}}\right)$ is not a rational multiple of $\pi$ by another short field-theoretic argument. [^31] If $\arccos\left(\frac{1}{\sqrt{2+\sqrt[3]{4}+\sqrt[5]{4}}}\right)$ is a rational multiple of $\pi$, then $K=\mathbb{Q}(\sqrt{2+\sqrt[3]{4}+\sqrt[5]{4}})$ is a subfield of a cyclotomic extension of $\mathbb{Q}$, but $K$ has $\mathbb{Q}(\sqrt[3]{4}+\sqrt[5]{4})$ as a subfield and the minimal polynomial of $\sqrt[3]{4}+\sqrt[5]{4}$ is $x^{15}-20x^{12}-12x^{10}+160x^{9}-1440x^{7}-640x^{6}+48x^{5}-8640x^{4}+1280x^{3}-1920x^{2}-3840x-1088$, so $\mathbb{Q}(\sqrt[3]{4}+\sqrt[5]{4})$ has the non-abelian Galois group $S_{3}\times\mathrm{Aut}(D_{5})$ [20]. Finally, $\langle 1,\sqrt[3]{2},\sqrt[5]{2}\rangle$ is not the fixed axis of any element of $F$, so we have that $\bar{x}\neq\alpha^{n}\cdot\bar{x}$ for any $n>0$ and $\bar{x}\in D$. Therefore the sets $(\alpha^{n}\cdot D)_{n\in\mathbb{N}}$ are pairwise disjoint.
@@
By A.3 and A.4, we have that the restriction of this equivalence relation to $S^{2}\setminus D$ is $\Delta^{0}_{3}$ (and therefore $\Delta^{1}_{2}$). The result then follows from $3.$ ∎
@@
Let $\tau$ and $\sigma$ be the two generators of $F_{2}$. The four sets $(A_{i})_{i\leq 4}$ given in [35, Thm. 4.2] are defined as follows:
@@
  • $x\in A_{4}$ if and only if the first letter in the reduced word of $x$ is $\sigma^{-1}$ and $x$ is not of the form $\sigma^{-n}$ for some $n$.
@@
This can almost certainly be improved to four pieces, the known optimal size of a paradoxical decomposition of $S^{2}$, by directly adapting the proof of [35, Thm. 4.5]. There is also no fundamental extra subtlety in writing out a $\Pi^{1}_{2}$ decomposition of the full unit ball, rather than just the unit sphere.